# Limits of Rauzy graphs of languages with subexponential complexity

### v1

@@ Abstract
To a subshift over a finite alphabet, one can naturally associate an infinite family of finite graphs, called its Rauzy graphs. We show that for a subshift of subexponential complexity the Rauzy graphs converge to the line $\mathbf{Z}$ in the sense of Benjamini-Schramm convergence if and only its complexity function $p(n)$ is unbounded and satisfies $\lim_{n}\frac{p(n+1)}{p(n)}=1$. We then apply this criterion to many examples of well-studied dynamical systems. If the subshift is moreover uniquely ergodic then we show that the limit of labelled Rauzy graphs if it exists can be identified with the unique invariant measure. In addition we consider an example of a non uniquely ergodic system recently studied by Cassaigne and Kaboré and identify a continuum of invariant measures with subsequential limits of labelled Rauzy graphs.
@@ 1 Introduction
Given an arbitrary sequence of finite graphs of uniformly bounded degree, I. Benjamini and O. Schramm [5] equipped it with a natural distributional limit which has since become a central object of investigation in many different contexts under the name of Benjamini-Schramm limit. The limiting object is not a graph, but rather a probability measure on the space of (isomorphism classes of) rooted graphs obtained by choosing uniformly at random a root in each of the graphs in the sequence and by considering all possible limits in the space of rooted graphs with local topology; see Section 2 for a precise definition. The key fact is that it is invariant under the equivalence relation of changing the root.
Applying this construction to a sequence of Rauzy graphs associated with a subshift, it is natural to ask how their Benjamini-Schramm limit is related to the original dynamical system and its invariant measures. In this paper we provide some answers to this question in the case of subshifts of subexponential complexity. In certain cases, the Benjamini–Schramm limits of Rauzy graphs give us an alternative description of the invariant measures for the original dynamical systems (see Section 4).
@@
For a factorial language ${\mathcal{L}}$, we denote by ${\vec{\operatorname{Rauz}}}(n)$ the $n$th Rauzy digraph of ${\mathcal{L}}$. That is the vertices of ${\vec{\operatorname{Rauz}}}(n)$ are the words of length $n$, while the arcs are words of length $n+1$. More precisely, for $w$ of length $n-1$, the word $awb$ is the arc from $aw$ to $wb$. Given two words $u$ and $v$ connected by an arc, the arc can be labelled by the last letter of $v$. We can thus speak about labelled Rauzy digraphs or, if we forget about the labelling of the edges, about unlabelled Rauzy digraphs. It is also possible to forget about orientation, replacing every arc by an undirected edge, and obtain labelled and unlabelled Rauzy graphs $\operatorname{Rauz}(n)$.
Our interest in the Benjamini-Schramm limits of families of Rauzy graphs initiated in the observation in [20] that the limit of De Brujin graphs $\{\mathcal{B}(k,N)\}_{N}$ (i.e., the Rauzy graphs of the full shift over the alphabet $\{0,...,k-1\}$) is the (delta-measure on the) Diestel-Leader graph $DL_{k,k}$ which is a Cayley graph of the lamplighter group $\mathbb{Z}_{k}\wr\mathbb{Z}$. Generalizing this result, the first-named author showed [27] that for languages coming from subshifts of finite type, the Benjamini-Schramm limit of Rauzy graphs is a measure supported on horocyclic products of trees that can be constructed from the subshift. On the other hand, Drummond studied the subshifts of linear complexity  [16] and proved that the Benjamini-Schramm limit of their unlabelled Rauzy digraphs is the (delta-measure on the) bi-infinite directed path $\vec{\mathbf{Z}}$ (see also [18, 31, 21] for related statements about Rauzy graphs of subshifts of linear complexity).
The main result of this paper is a characterisation of the languages whose Rauzy (di)graphs ${\vec{\operatorname{Rauz}}}(n)$ converge to $\vec{\mathbf{Z}}$:
@@
Let ${\mathcal{L}}$ be a factorial and almost prolongable language with complexity function $p(n)$. The following are equivalent:
@@
    The Rauzy graphs $\operatorname{Rauz}(n)$ converge to the line $\mathbf{Z}$,
@@
    $p(n)$ is unbounded, grows subexponentially and the limit of the Rauzy graphs $\operatorname{Rauz}(n)$ exists,
@@
    $p(n)$ is unbounded, grows subexponentially and the limit of the Rauzy digraphs ${\vec{\operatorname{Rauz}}}(n)$ exists,
@@
    $p(n)$ is unbounded and $\lim_{n}\frac{p(n+1)}{p(n)}=1$.
Theorem 1.1 directly applies to languages coming from bi-sided subshifts, as they are both factorial and prolongable, hence, almost prolongable. While the language of a one-sided subshift is in general not almost prolongable, it is the case for a broad class of subshifts. Surjective subshifts, in particular minimal ones, have prolongable language of subwords. Moreover, if a subshift has a dense orbit or, more generally, if it has a finite set with dense orbit then the language is almost prolongable (see Subsection 2.3). Therefore, we obtain the following corollary of Theorem 1.1:
Corollary 1.2.
Let $\Sigma$ be a bi-sided subshift or a one-sided subshift. For one-sided subshifts, suppose moreover that $\Sigma$ has a finite subset with a dense orbit. The following are equivalent:
  1. 1.
    The Rauzy graphs $\operatorname{Rauz}(n)$ converge to the line $\mathbf{Z}$,
  2. 2.
  3. 3.
    $p(n)$ is unbounded and $\lim_{n}\frac{p(n+1)}{p(n)}=1$.
@@
    $p(n)\sim n^{\gamma}$ for some real number $\gamma$,
@@
In Section 4, we will see that Corollary 1.2 and Corollary 1.3 apply to many interesting dynamical systems of different origins, including interval exchange transformations and interval exchange transformations with flips, interval translation maps, polygonal and polyhedral billiards, Toeplitz subshifts, multidimensional continuous fractions and more.
Since Rauzy digraphs come with a natural labelling, we are also interested in the convergence of the labelled Rauzy digraphs. If the limit exists then Theorem 1.1 implies that it will be a probability measure on rooted labelled digraphs supported on $\vec{\mathbf{Z}}$ and invariant under the change of the root and so can be thought of as a probability measure on the set of bi-infinite words in the alphabet $\mathcal{A}$ invariant under the shift. It is therefore not surprising that unique ergodicity is a natural condition sufficient for the existence of the Benjamini-Schramm limit of labelled Rauzy digraphs.
@@
Section 2 contains all the necessary definitions and Theorem 2.6 which generalizes Theorem 1.1 and implies Corollary 1.2 for bi-sided subshifts. It also contains the proof of Corollary 1.3, which is split between Corollaries 2.7 and 2.13 for bi-sided subshifts. The statements about one-sided subshifts in Corollary 1.2 and Corollary 1.3 are proved with the help of an additional technical result: Proposition 2.14. Section 3 deals with the limits of labelled Rauzy graphs. Section 4 is devoted to applications of the above to various concrete examples of dynamical systems. Finally, Section 5 contains an example of a non-uniquely ergodic subshift where we identify subsequential limits of labelled Rauzy graphs and show that they give us a continuum of invariant measures.
@@ 2 Limits of Rauzy graphs of languages with subexponential complexity
A rooted graph is a connected graph with a specified vertex, which is called the root. Let us denote by $\mathcal{G}$ the space of rooted graphs of uniformly bounded degree, where two rooted digraphs are close if they coincide on a big ball around their roots. One defines similarly the spaces of rooted digraphs, of rooted labelled graphs and of rooted labelled digraphs. We will write $\mathcal{G}$ as a placeholder for any of the four above spaces. The topology of $\mathcal{G}$ is induced by the metric: $d((G_{1},v_{1}),(G_{2},v_{2}))=2^{-r}$ where $r$ is the biggest radius such that $G_{1}$ and $G_{2}$ have isomorphic balls of radius $r$ around their respective root. The metric is complete and separable. Moreover, the space $\mathcal{G}$ of rooted (di)graphs (labelled or not) of uniformly bounded degree is compact. Since here we study limits of Rauzy graphs of languages over an alphabet with $k$ letters, all our graphs are of degree bounded by $k$. Since $\mathcal{G}$ is separable and complete metrizable, the space $\mathbf{P}(\mathcal{G})$ of Borel probability measures on $\mathcal{G}$ is compact in the weak-*** topology. For a sequence $(G_{i})_{i}$ of finite graphs in $\mathcal{G}$ let us choose the root in each of them uniformly at random. This means that we define, for each $i$, the measure
@@
An important property of the Benjamini-Schramm limit is its unimodularity, that is, invariance under the change of the root (see [1, 24] for details). A specific case of convergence of random (di)graphs is when they converge to a measure $\mu$ supported on one rooted (di)graph $(G,v)$ with $G$ vertex-transitive. In that case, we say that the random (di)graphs converge to the graph $G$.
@@ 2.1 Rauzy (di)graphs converging to the (directed) line
(1$\implies$(2). It is clear that the convergence of the Rauzy digraphs implies the convergence of the Rauzy graphs. It hence remains to show that it also implies that ${\mathcal{L}}$ is almost prolongable. It follows from the definition that for *$*\in\{l,r\}$ we have
@@
But then, using that the convergence of ${\vec{\operatorname{Rauz}}}(n)$ to $\vec{\mathbf{Z}}$ implies $\lim_{n}\frac{r(n)}{p(n)}=1$, the above inequality implies that $\lim_{n}\frac{s_{*}(n)}{p(n)}=\lim_{n}\frac{e_{*}(n)}{p(n)}=0$ for *$*\in\{l,r\}$. In particular, ${\mathcal{L}}$ is almost prolongable. ∎
@@
Let ${\mathcal{L}}$ be a factorial language with complexity $p(n)$. Then the following are equivalent:
@@
    ${\mathcal{L}}$ is almost prolongable and the Rauzy graphs $\operatorname{Rauz}(n)$ converge to the line $\mathbf{Z}$,
@@
    ${\mathcal{L}}$ is almost prolongable, $p(n)$ is unbounded, grows subexponentially, the limit of the Rauzy digraphs ${\vec{\operatorname{Rauz}}}(n)$ exists,
@@
    ${\mathcal{L}}$ is almost prolongable, $p(n)$ is unbounded, grows subexponentially, the limit of the Rauzy graphs $\operatorname{Rauz}(n)$ exists,
@@
    ${\mathcal{L}}$ is almost prolongable, $p(n)$ is unbounded and $\lim_{n}\frac{p(n+1)}{p(n)}=1$.
@@
(3$\implies$(4) The limit of the $\operatorname{Rauz}(n)$ is the underlying graph of the limit of the ${\vec{\operatorname{Rauz}}}(n)$.
@@
(1$\iff$(2$\iff$(5) By a theorem of Ehrenfeucht and Rozenberg [17], for a factorial language ${\mathcal{L}}$ either $p(n)$ is bounded or $p(n)\geq n+1$ for every $n$. In particular, the condition $p(n)$ is unbounded is equivalent to $\lim_{n}p(n)=\infty$. Thus, by Lemma 2.4, it is enough to show that, for a factorial language, we have $\lim_{n}\frac{r(n)}{p(n)}=1$ if and only if ${\mathcal{L}}$ is almost prolongable and $\lim_{n}\frac{p(n+1)}{p(n)}=1$. Observe that the last equality is equivalent to $\lim_{n}\frac{p(n+1)-p(n)}{p(n)}=0$.
@@
(1$\implies$(3) We have to show that Condition (1) implies that ${\mathcal{L}}$ is almost prolongable, and $p(n)$ is unbounded and grows subexponentially. The first part follows from (2) and the second part from (5).
@@
If the complexity function $p(n)$ of a regular factorial language is subexponential then is satisfies $\lim_{n\to\infty}\frac{p(n+1)-p(n)}{p(n)}=0.$
@@
It is known that a regular language has either exponential or polynomial complexity (see [33]). Moreover, in the latter case, the complexity function is a quasi-polynomial and, hence, satisfies $\lim_{n\to\infty}\frac{p(n+1)-p(n)}{p(n)}=0$. ∎
@@
The Rauzy graphs of any regular factorial almost prolongable language converge in Benjamini-Schramm sense to $\mathbf{Z}$.

### v2

@@ Abstract
To a subshift over a finite alphabet, one can naturally associate an infinite family of finite graphs, called its Rauzy graphs. We show that for a subshift of unbounded subexponential complexity the Rauzy graphs converge in the sense of Benjamini-Schramm if and only if their empirical spectral measures converge, which is equivalent to the complexity function $p(n)$ to satisfy $\lim_{n}\frac{p(n+1)}{p(n)}=1$. The limit in this case is the infinite line $\mathbf{Z}$. If the subshift is moreover uniquely ergodic, the limit of labelled Rauzy digraphs if it exists can be identified with the unique invariant measure. We also consider a non uniquely ergodic example recently studied by Cassaigne and Kaboré and identify a continuum of invariant measures with subsequential limits of labelled Rauzy digraphs.
@@ 1 Introduction
Given an arbitrary sequence of finite graphs of uniformly bounded degree, I. Benjamini and O. Schramm [7] equipped it with a natural distributional limit which has since become a central object of investigation in many different contexts under the name of Benjamini-Schramm limit. The limiting object is not a graph, but rather a probability measure on the space of (isomorphism classes of) rooted graphs obtained by choosing uniformly at random a root in each of the graphs in the sequence and by considering all possible limits in the space of rooted graphs with local topology. The key fact is that it is unimodular; see Section 2 for a precise definition. In this paper, when talking about convergence of graphs, we will always mean convergence in the sense of Benjamini-Schramm. In the special case when the limiting measure of a sequence of finite graphs $(G_{i})_{i}$ is supported on one (unimodular, hence transitive) graph $G$ we will say that $(G_{i})_{i}$ converge to $G$.
Applying this construction to a sequence of Rauzy graphs associated with a subshift, it is natural to ask how their Benjamini-Schramm limit is related to the original dynamical system and its invariant measures. In this paper we provide some answers to this question in the case of subshifts of subexponential complexity. It turns out that in this case convergence of graphs can be detected by spectral convergence. In various examples, the Benjamini–Schramm limits of Rauzy graphs give us an alternative description of the invariant measures for the original dynamical systems (see Section 4).
@@
For a factorial language ${\mathcal{L}}$, we denote by ${\vec{\operatorname{Rauz}}}(n)$ the $n$th Rauzy digraph of ${\mathcal{L}}$. That is, the vertices of ${\vec{\operatorname{Rauz}}}(n)$ are the words of length $n$, while the arcs are words of length $n+1$. More precisely, for $w$ of length $n-1$, the word $awb$ is the arc from $aw$ to $wb$. Given two words $u$ and $v$ connected by an arc, the arc can be labelled by the last letter of $v$. We can thus speak about labelled Rauzy digraphs or, if we forget about the labelling of the arcs, about unlabelled Rauzy digraphs. It is also possible to forget about orientation, replacing every arc by an undirected edge, and obtain labelled and unlabelled Rauzy graphs $\operatorname{Rauz}(n)$.
Our interest in the Benjamini-Schramm limits of families of Rauzy graphs initiated in the observation in [22] that the limit of De Brujin graphs $\{\mathcal{B}(k,N)\}_{N}$ (i.e., the Rauzy graphs of the full shift over the alphabet $\{0,...,k-1\}$) is the (delta-measure on the) Diestel-Leader graph $DL_{k,k}$ which is a Cayley graph of the lamplighter group $\mathbf{Z}_{k}\wr\mathbf{Z}$. Generalizing this result, the first-named author showed [29] that for languages coming from subshifts of finite type, the Benjamini-Schramm limit of Rauzy graphs is a measure supported on horocyclic products of trees that can be constructed from the subshift. On the other hand, Drummond studied the subshifts of linear complexity  [18] and proved that the Benjamini-Schramm limit of their unlabelled Rauzy digraphs is the (delta-measure on the) bi-infinite directed path $\vec{\mathbf{Z}}$ (see also [20, 34, 23] for related statements about Rauzy graphs of subshifts of linear complexity).
The question about the Benjamini-Schramm limit of de Brujin graphs was motivated by spectral computations that appeared in [5]. Indeed, if a sequence of finite graphs converges, then the associated empirical spectral measures (aka counting measures on the spectra of the corresponding adjacency matrices) converge to the expected spectral measure on their Benjamini-Schramm limit, see Section 2 for details. Spectral convergence however does not in general imply graph convergence.
The main result of this paper is a characterisation of convergence, spectral and in the Benjamini-Schramm sense, of Rauzy graphs of languages with subexponential complexity.
@@
Let ${\mathcal{L}}$ be a factorial and almost prolongable language with unbounded complexity function $p(n)$. The following are equivalent:
@@
    $p(n)$ grows subexponentially and the limit of the Rauzy graphs $\operatorname{Rauz}(n)$ exists.
@@
    The Rauzy graphs $\operatorname{Rauz}(n)$ converge to the line $\mathbf{Z}$.
@@
    $p(n)$ grows subexponentially and the empirical spectral measures $\mu_{\operatorname{Rauz}(n)}$ converge.
@@
    The empirical spectral measures $\mu_{\operatorname{Rauz}(n)}$ converge to the measure that has density $\frac{1}{\pi\sqrt{4-x^{2}}}dx$ with respect to the Lebesgue measure on $[-2,2]$.
@@
    $p(n)$ satisfies $\lim_{n}\frac{p(n+1)}{p(n)}=1$.
  6. 6.
    $p(n)$ grows subexponentially and the limit of the Rauzy digraphs ${\vec{\operatorname{Rauz}}}(n)$ exists.
  7. 7.
    The Rauzy digraphs ${\vec{\operatorname{Rauz}}}(n)$ converge to the directed line $\vec{\mathbf{Z}}$.
Theorem 1.1 directly applies to languages coming from bi-sided subshifts, as they are both factorial and prolongable, hence, almost prolongable. While the language of a one-sided subshift is in general not almost prolongable, it is the case for a broad class of subshifts. Surjective subshifts, in particular minimal ones, have prolongable language of subwords. Moreover, if a subshift has a dense orbit or, more generally, if it has a finite set with dense orbit then the language is almost prolongable (see Subsection 2.3).
@@
    $p(n)\sim n^{\gamma}$ for some real number $\gamma$.
@@
In Section 4, we will see that Theorem 1.1 and Corollary 1.2 apply to many interesting dynamical systems of different origins, including interval exchange transformations and interval exchange transformations with flips, interval translation maps, polygonal and polyhedral billiards, Toeplitz subshifts, multidimensional continuous fractions and more.
Since Rauzy digraphs come with a natural labelling, we are also interested in the convergence of the labelled Rauzy digraphs. If the limit exists, then Theorem 1.1 implies that it will be a probability measure on rooted labelled digraphs supported on $\vec{\mathbf{Z}}$. It can be thought of as a probability measure on the set of bi-infinite words in the alphabet $\mathcal{A}$ invariant under the shift (as a consequence of unimodularity, see Section 2 for details). It is therefore not surprising that unique ergodicity is a natural condition sufficient for the existence of the Benjamini-Schramm limit of labelled Rauzy digraphs.
@@
Section 2 contains all the necessary definitions and the proof of Theorem 1.1. It also contains the proof of Corollary 1.2, which is split between Corollaries 2.10 and 2.16 for bi-sided subshifts. The statements about one-sided subshifts to which we can apply Theorem 1.1 and Corollary 1.2 are proved with the help of an additional technical result: Proposition 2.17. Section 3 deals with the limits of labelled Rauzy graphs. Section 4 is devoted to applications of the above to various concrete examples of dynamical systems. Finally, Section 5 contains an example of a non-uniquely ergodic subshift where we identify subsequential limits of labelled Rauzy graphs and show that they give us a continuum of invariant measures.
@@ 2 Limits of Rauzy graphs of languages with subexponential complexity
A rooted graph is a connected graph with a specified vertex, which is called the root.
Let us denote by $\mathcal{G}$ the space of (isomorphism classes of) rooted graphs of uniformly bounded degree, where two rooted graphs are close if they coincide on a big ball around their roots. One defines similarly the spaces of rooted digraphs, of rooted labelled graphs and of rooted labelled digraphs. We will write $\mathcal{G}$ as a placeholder for any of the four above spaces. The topology of $\mathcal{G}$ is induced by the metric: $d((G_{1},v_{1}),(G_{2},v_{2}))=2^{-r}$ where $r$ is the biggest radius such that $G_{1}$ and $G_{2}$ have isomorphic balls of radius $r$ around their respective root. The metric is complete and separable. Moreover, the space $\mathcal{G}$ of rooted (di)graphs (labelled or not) of uniformly bounded degree is compact. Since here we study limits of Rauzy graphs of languages over an alphabet with $k$ letters, all our graphs are of degree bounded by $k$. Since $\mathcal{G}$ is separable and complete metrizable, the space $\mathbf{P}(\mathcal{G})$ of Borel probability measures on $\mathcal{G}$ is compact in the weak topology.
For a sequence $(G_{i})_{i}$ of finite graphs in $\mathcal{G}$ let us choose the root in each of them uniformly at random. This means that we define, for each $i$, the measure
@@
An important property of the Benjamini-Schramm limit is its unimodularity that we will now briefly define (see [2] for details).
Definition 2.3.
A flagged rooted graph is a rooted graph $(G,v)$ with a chosen edge from $E(G)$ incident to the root and oriented such that $v$ is its tail. Consider the space of flagged rooted graphs of uniformly bounded degree $\mathcal{FG}$. There exists a natural map $\pi:\mathcal{FG}\to\mathcal{G}$ that simply forgets the flag. Given a random rooted graph $\mu\in\mathbf{P}(\mathcal{G})$, we consider a measure $\mu^{f}$ on $\mathcal{FG}$ given by assigning to each flagged rooted graph the mass of the underlying rooted graph. Consider the operator $R$ on the space $\mathcal{FG}$ acting by replacing the root to the head of the flag, and reversing the orientation of the flag. We say that random rooted graph $\mu$ is unimodular if the measure $\mu^{f}$ is invariant under $R$. The same definition applies to digraphs and to labelled graphs and labelled digraphs. There one has to choose a flag among all the edges incident to the root (independently of their orientation and labels).
A specific case of Benjamini-Schramm convergence is when the limit is a measure $\mu$ supported on one rooted (di)graph $(G,v)$ with $G$ unimodular and hence vertex-transitive. As Theorem 1.1 shows, this is the case for Rauzy graphs of languages of subexponential complexity, as their limit, if exists, is (the delta-measure on) the infinite line $\mathbf{Z}$. In this particular case the limit of labelled Rauzy digraphs becomes a measure on the set $\mathcal{A}^{\mathbf{Z}}$ of arc-labellings of the oriented line with letters from $\mathcal{A}$, and we have the following Lemma.
Lemma 2.4.
Let $\mu$ be a unimodular measure on the space of labelled digraphs supported on $A^{\mathbf{Z}}$ that we realize as the set of arc-labellings of the oriented infinite line $\vec{\mathbf{Z}}$. Then $\mu$ is shift invariant.
Proof.
The measure $\mu^{f}$ is supported on two copies of $A^{\mathbf{Z}}$, that is, on the set $A^{\mathbf{Z}}\times\{+1,-1\}$, where $+1$ corresponds to the flag pointing to the right and $-1$ corresponds to the flag pointing to the left. Moreover, the measure $\mu^{f}$ is exactly the product $\mu\otimes\{\delta_{+1},\delta_{-1}\}$. Let $T$ be the shift on $A^{\mathbf{Z}}$. In terms of $T$, the operator $R$ can be expressed as follows
$R(x,\sigma)=(T^{\sigma}x,-\sigma),\quad x\in A^{\mathbf{Z}},\ \sigma\in\{{+1},{-1}\}.$
Let $B$ be a measurable subset of $A^{\mathbf{Z}}$. Then due to $R$-invariance of $\mu^{f}$ we have
$\mu(T(B))=\mu^{f}(T(B)\times\{-1\})=\mu^{f}(B\times\{+1\})=\mu(B).$
Therefore, the measure $\mu$ is invariant under $T$. ∎
Given a finite graph $G$, let $A_{G}$ be its adjacency matrix and $\sigma(A_{G})\subset\mathbf{R}$ be the set of eigenvalues of $A_{G}$. Let $m_{\lambda}$ be the multiplicity of an eigenvalue $\lambda\in\sigma(A_{G})$. The empirical spectral measure $\mu_{G}$ of $G$ is the following measure
$\mu_{G}=\frac{1}{\lvert V(G)\rvert}\sum_{\lambda\in\sigma(A_{G})}m_{\lambda}\delta_{\lambda},$
where $\delta_{\lambda}$ is the Dirac measure at point $\lambda$. It can be shown that for a sequence of finite graphs $(G_{i})_{i}$ converging in the Benjamini-Schramm sense to a random rooted graph $\nu\in\mathbf{P}(\mathcal{G})$, the empirical spectral measures $\mu_{G_{i}}$ converge to the expected spectral measure of $\nu$ which is the average of spectral measures $\mu_{G,v}$ of a graph $G$ computed at the vertex $v$, where the rooted graph $(G,v)$ belongs to the support of $\nu$; see, e.g., [31] for spectral measures on infinite graphs and [1] for expected spectral measures on random rooted graphs. (Note that expected spectral measure can be defined for any random rooted graphs and for a finite graph with the uniform root, the expected spectral measure and the empirical spectral measure coincide.) In general, convergence of spectral measures $\mu_{G_{i}}$, for a sequence of finite graphs $(G_{i})_{i}$, does not imply convergence of $(G_{i})_{i}$.
As follows from 1.1, the only random rooted graph that we have to deal with in this paper is the delta-measure on a transitive graph, namely, the infinite line $\mathbf{Z}$, in which case there is no need to take the average and the limit is just the spectral measure of the adjacency operator on $\mathbf{Z}$ which is independent on the vertex $v\in\mathbf{Z}$ and which can be easily found, see, e.g., [24][Prop.3.2].
@@ 2.1 Rauzy (di)graphs converging to the (directed) line
(1$\implies$(2). It is clear that the convergence of the Rauzy digraphs implies the convergence of the Rauzy graphs. It hence remains to show that it also implies that ${\mathcal{L}}$ is almost prolongable. It follows from the definition that for $*\in\{l,r\}$ we have
@@
But then, using that the convergence of ${\vec{\operatorname{Rauz}}}(n)$ to $\vec{\mathbf{Z}}$ implies $\lim_{n}\frac{r(n)}{p(n)}=1$, the above inequality implies that $\lim_{n}\frac{s_{*}(n)}{p(n)}=\lim_{n}\frac{e_{*}(n)}{p(n)}=0$ for $*\in\{l,r\}$. In particular, ${\mathcal{L}}$ is almost prolongable. ∎
@@
Let ${\mathcal{L}}$ be a factorial and almost prolongable language with unbounded complexity $p(n)$. Then the following are equivalent:
@@
    The Rauzy digraphs ${\vec{\operatorname{Rauz}}}(n)$ converge to the directed line $\vec{\mathbf{Z}}$.
@@
    The Rauzy graphs $\operatorname{Rauz}(n)$ converge to the line $\mathbf{Z}$.
@@
    $p(n)$ grows subexponentially and the limit of the Rauzy digraphs ${\vec{\operatorname{Rauz}}}(n)$ exists.
@@
    $p(n)$ grows subexponentially and the limit of the Rauzy graphs $\operatorname{Rauz}(n)$ exists.
@@
    $p(n)$ satisfies $\lim_{n}\frac{p(n+1)}{p(n)}=1$.
  6. 6.
    The empirical spectral measures $\mu_{\operatorname{Rauz}(n)}$ converge to the measure that has density $\frac{1}{\pi\sqrt{4-x^{2}}}dx$ with respect to the Lebesgue measure on $[-2,2]$.
  7. 7.
    $p(n)$ grows subexponentially and the empirical spectral measures $\mu_{\operatorname{Rauz}(n)}$ converge.
@@
(1$\implies$(2) and (3$\implies$(4) The limit of the $\operatorname{Rauz}(n)$ is the underlying graph of the limit of the ${\vec{\operatorname{Rauz}}}(n)$.
@@
(1$\iff$(2$\iff$(5) We will in fact prove something slightly stronger. Namely, that for a general language ${\mathcal{L}}$ we have (1$\iff$[(2) and ${\mathcal{L}}$ is almost prolongable]$\iff$[(5), ${\mathcal{L}}$ is almost prolongable and $p(n)$ is unbounded]. By a theorem of Ehrenfeucht and Rozenberg [19], for a factorial language ${\mathcal{L}}$ either $p(n)$ is bounded or $p(n)\geq n+1$ for every $n$. In particular, the condition $p(n)$ is unbounded is equivalent to $\lim_{n}p(n)=\infty$. Thus, by Lemma 2.6, it is enough to show that, for a factorial language, we have $\lim_{n}\frac{r(n)}{p(n)}=1$ if and only if ${\mathcal{L}}$ is almost prolongable and $\lim_{n}\frac{p(n+1)}{p(n)}=1$. Observe that the last equality is equivalent to $\lim_{n}\frac{p(n+1)-p(n)}{p(n)}=0$.
@@
(1$\implies$(3) Once again we prove something sligthly stronger than announced: for a general language ${\mathcal{L}}$ we have (1$\implies$[(3), ${\mathcal{L}}$ is almost prolongable and $p(n)$ is unbounded]. We have to show that Condition (1) implies ${\mathcal{L}}$ is almost prolongable, and $p(n)$ is unbounded and grows subexponentially. The first part follows from (2) and the second part from (5).
(7$\implies$(5) Convergence of empirical spectral measures $\mu_{\operatorname{Rauz}(n)}$ implies the convergence of their positive moments. The second moment of the empirical measure of a simple graph $(V,E)$ is double the average degree which is $\frac{2\lvert E\rvert}{\lvert V\rvert}$. In general case when the graph might have loops and double edges, the second moment of the empirical measure is
$\frac{1}{\lvert V\rvert}\sum_{v\in V}\lvert\{(e_{1},e_{2})\in E\colon e_{1}\textnormal{ and }e_{2}\textnormal{ have the same endpoints, one of which is }v\}\rvert$
$\frac{1}{\lvert V\rvert}\big{(}\sum_{v\in V}\lvert\{\textnormal{loops in }v\}\rvert^{2}+\sum_{v,u\in V,\ u\not=v}\lvert\{\textnormal{edges between }u\textnormal{ and }v\}\rvert^{2}\big{)}$
In the case of Rauzy graphs, we have no more than $\lvert{\mathcal{A}}\rvert^{2}$ double edges and no more than $\lvert{\mathcal{A}}\rvert$ loops. Hence
[…]
Hence, the limit $\lim_{n}\frac{p(n+1)}{p(n)}$ exists. Due to subexponential complexity of the subshift, the only possible value of this limit is $1$.
(6$\implies$(7) In case we know that the empirical measures $\mu_{\operatorname{Rauz}(n)}$ converge to the measure that has density $\frac{1}{\pi\sqrt{4-x^{2}}}dx$, then it follows from the Equation 7 that the complexity is subexponential.
(2$\implies$(6) is done in [1] in the general case and can easily be verified in our particular case when the limit is (a delta-measure on) a unique transitive graph, $\mathbf{Z}$. ∎
Remark 2.9.
As seen in the proof of Theorem 2.8, for a general language ${\mathcal{L}}$, the convergence of the Rauzy digraphs ${\vec{\operatorname{Rauz}}}(n)$ to the directed line $\vec{\mathbf{Z}}$ implies both that ${\mathcal{L}}$ is almost prolongable and $p(n)$ is unbounded. Similarly, Conditions 2 and 6 of Theorem 2.8 each implies that $p(n)$ is unbounded.
@@
If the complexity function $p(n)$ of a regular factorial language is subexponential then is satisfies $\lim_{n\to\infty}\frac{p(n+1)}{p(n)}=1.$
@@
It is known that a regular language has either exponential or polynomial complexity (see [36]). Moreover, in the latter case, the complexity function is a quasi-polynomial and, hence, satisfies $\lim_{n\to\infty}\frac{p(n+1)}{p(n)}=1$. ∎
@@
The Rauzy graphs of any regular factorial almost prolongable language converge in the Benjamini-Schramm sense to $\mathbf{Z}$.